# Whispering in the dark

### v2

@@ 3.3 X-ray luminosity from a Keplerian accretion disc
where is the wind loss rate of the OB star, is the orbital separation, is the accretion radius of the BH, is the wind velocity of the O star, is the relative velocity of the BH with respect to the wind velocity (see Eqs. (1)-(8) of Sen et al. 2021), and represents the self-limiting effect of the Eddington luminosity,
@@
where , , and are the mass, radius, and Eddington factor of the OB star companion.
3.4 X-ray luminosity without an accretion disc
When the specific angular momentum carried by the OB star wind matter is insufficient to form a Keplerian accretion disc around the BH, the wind matter can spiral into the BH following the magnetic field lines. During this in-fall, we show in Sect. 2.2.3 that electrons can receive a portion of the viscously dissipated energy of the in-falling matter, and emit radiation in the X-ray band. The X-ray luminosity is
where is the radiative efficiency and is the net mass-accretion rate at the radius of the ISCO of the BH.
Recent simulations of hot accretion flows have revealed the presence of outflows in ADAFs, such that the mass inflow rate of the accreting gas decreases as the material approaches the event horizon of the BH (Stone et al. 1999; Yuan & Bu 2010). We assume the mass-accretion rate scales with the radius from the BH by a power-law (, Stone et al. 1999; Igumenshchev et al. 2000; Stone & Pringle 2001; Yuan & Bu 2010). Hence, the relation between the net mass-accretion rate at the radius of the innermost stable circular orbit and the mass-accretion rate at the accretion radius is given by
@@ 3.4 X-ray luminosity without an accretion disc
Xie & Yuan (2012, Fig. 1) studied the radiative efficiency of hot accretion flows around a stellar-mass BH orbiting an OB star companion. They derive as a function of the net mass-accretion rate (i.e. = (,)), for three different values of the viscous dissipation parameter (defined in Sect. 2.2.3). From their figure 1, we extract the values of the radiative efficiency as a function of the net mass-accretion rate to calculate the X-ray luminosity during the BH+OB star phase in our models for viscous dissipation parameter . While we show in Sect. 2 that viscous heating of electrons can be efficient even in the case of stellar mass BHs with OB star companions, we present our results for all three values above to derive a lower limit on the number of faint X-ray sources from BH+OB star binaries for the most inefficient case of viscous heating.
@@ 4 Results
4.1 Mass-accretion rate
Figure 1 shows the distribution of the mass-accretion rate at the accretion radius (left panel) and the net mass-accretion rate near the event horizon of the BH (right panel). Higher mass-accretion rates are reached for shorter orbital period systems (Eq. (22), expressed in terms of orbital separation). The peak in the mass-accretion rate distribution is and for 1.4-10 d and 10-100 d initial orbital periods respectively. The shortest-period systems can also form an accretion disc around the BH (Sen et al. 2021), denoted by the black hatching. We note however that there is a significant contribution from the 10-100 d binaries to the population of BH+OB star systems that can form an accretion disc (see Sect. 4.3).
The highest mass-accretion rates () is mostly from models in the 10-100 d initial orbital period range. The shortest initial orbital period models ( d) enter a contact phase and merge on the main sequence (Menon et al. 2021). Many of the short-period binary models (2-10 d) in our grid do not survive their prior Case A and Case AB mass transfer phase (Sen et al. 2022, Fig. 1 of, Pauli et al. 2022, Appendix C of) to reach the BH+OB star phase.
The total number of short-period (1.4-10 d) binary models contributing to the distribution of mass-accretion rate is much smaller than the number of models with longer initial orbital periods (10-100 d). Correspondingly, the peak in the distribution of mass-accretion rate does not occur at the highest mass-accretion rates but at lower mass-accretion rates where more models contribute to the histogram function (Eq. (33)), despite our assumed intrinsic period distribution favouring short-period binaries (Eq. (34)).
The mass-accretion rate typically decreases for models with longer orbital periods (Eq. (22)). The drop in the number of systems with mass-accretion rates below is produced by our assumption that longer-period binaries are less likely to be born (Eq. (34)). Hence, the peak in the distribution of mass-accretion rates arises at 10-12-10-11 yr-1, from BH+OB star binaries with intermediate orbital periods.
We assume that mass outflows reduce the mass-accretion rate from the accretion radius to the BH event horizon (Eq. 32). The right panel shows that the peak in the distribution of the net mass-accretion rate near the event horizon of the BH is 2 orders of magnitude lower than that at the accretion radius. This reduction in mass-accretion rate results in a proportionate decrease in X-ray luminosity (Eq. 31) predicted from ADAFs around BHs in the BH+OB star binaries. We note that observations have confirmed the presence of outflows in Sgr A* (Hawley & Balbus 2002; Yuan et al. 2003; Igumenshchev et al. 2003). In case the same may not true for accretion onto stellar-mass BHs (e.g. see Fender & Gallo 2014), our predictions for the X-ray luminosity will be 2 orders of magnitude higher than presented in the next section.
@@ 4.2 X-ray luminosity
Figure 2 shows the X-ray luminosity distribution from BH+OB star binaries according to our binary evolution models due to accretion onto the BH (coloured histograms) and shocks in the wind of the OB star companion (black step histogram). In all the panels, the X-ray luminosity in the range erg s-1 originate solely from a Keplerian accretion disc (Shakura 1973) around the BHs (Eq. 25). These strong X-ray sources comprise 7.85% of the entire population of BH+OB star binaries in our grid. The orbital period distribution (see figure 6 of Langer et al. 2020) of the BH+OB star binaries peaks above 100 d, where the strong wind velocity of the OB star companion disfavours the formation of an accretion disc (Sen et al. 2021). Hence, most BH+OB star binaries in our model grid do not form an accretion disc. We call the BH+OB star binaries with X-ray luminosity above erg s-1 ‘X-ray-bright’.
We find that the contribution from binary models in the 40-90 range to the population of X-ray-bright BH+OB star binaries is negligible. We identify two reasons. Firstly, the IMF strongly disfavours the most massive binaries in the population. Secondly, the most massive models enter a contact phase up to an initial orbital period of 3 d (figures C1-C3 of Pauli et al. 2022). Hence, the shortest-period models, likely to form an accretion disc during a BH+OB phase, do not reach their BH+OB star phase.
The X-ray luminosity distribution shows a second, broader peak at luminosities in the range from erg s-1 which is associated with systems having ADAFs (Narayan & Yi 1995) around BHs that do not have Keplerian accretion disc around them (Sect. 3.4). The largest contribution comes from the binaries with initial primary masses of 20-40 . A large fraction of the primaries in the initial mass range of 10-20 have helium core mass smaller than 6.6 such that they are not expected to collapse into BHs while the contribution from the 40-90 range is suppressed by the IMF and comparatively shorter lifetimes of more massive O-star companions.
@@
The X-ray luminosity of observed Be X-ray binaries is higher than 1034 erg s-1, even for the faint persistent sources (Cheng et al. 2014). For inefficient viscous heating of electrons, the X-ray luminosity predicted from the BH+O star binaries remain below the lower limit of the X-ray luminosity from Be X-ray binaries (top and middle panels). For = 0.5 (bottom panel), only 2% of the BH+O star binaries may have X-ray luminosity between erg s-1, which is lower than the fraction of BH+O star binaries predicted to have a Keplerian accretion disc around the BH. We discuss this in more detail in Sect. 6.2.
@@ 4.3 X-ray-bright systems
The left panel of Fig. 3 shows the distribution of the Roche lobe filling factor of the OB star companion when an accretion disc can form during the BH+OB star phase of the binary models. We find a distinct bimodal distribution, partly as a natural consequence of the binary orbital periods (1.4 d-10d vs 10-100 d) that contribute to the different peaks (0.3 vs 0.9). The filling factors of models arising from 1.4-10 d models peak at 0.9. The orbital periods are short enough that the accretion disc can form only towards the end of the main sequence of the O star companions when the stellar radius increases and the wind speed decreases. On the other hand, the contribution from 10-100 d binary models comes primarily from BHs with B star companions, where the terminal velocity of the B star wind is a factor of two lower than O stars. Moreover, many binary models with small initial mass ratios merge on the main sequence at the 1.4-10 d orbital period range (e.g. figure 1 and figure F1 of Sen et al. 2022).
@@ 4.4 X-ray-faint systems
Figure 4 shows the orbital period and OB star luminosity of the BH+OB star binaries that are predicted to emit X-ray emission above 1031 erg s-1, the observational threshold of current X-ray telescopes/surveys (Crowther et al. 2022). For = 0.001, the predicted distribution of orbital periods peak near 10 d. The mass-accretion rate falls inversely with the square of the orbital separation (Eq. 22), thereby making higher orbital period binaries less likely to overcome the X-ray luminosity threshold in the LMC. At the shortest orbital periods, all our binary models do not survive the mass transfer phases before the BH+OB star phase (Sect. 4.1), due to which the probability distribution decreases below 10 d.
The wind-mass-loss rate of the OB stars increases with luminosity (see e.g. Kudritzki & Puls 2000; Puls et al. 2008) and mass (Vink et al. 2001a). BHs in orbit with more massive (luminous) stars at the same orbital period have a greater mass-accretion rate than less massive (luminous) stars (Eq. 22), as the mass accretion rate is proportional to the wind mass-loss rate. Consequently, the luminosity of the OB star has to be higher at long orbital periods to keep the mass-accretion rate onto the BH high enough to produce observable X-ray emission. As such, we see that the distribution of observable BH+OB star binaries extends towards higher luminosities at higher orbital periods.
For = 0.5 (right panel of Fig. 4), the radiative efficiency increases by 2 orders of magnitude compared to = 0.001, which leads to higher X-ray luminosity produced from the same mass-accretion rate. We see a larger contribution from longer-period binaries to the distribution of BH+OB star binaries with observable X-ray emission. The increased radiative efficiency compensates for the lower mass-accretion rate at longer periods. For high efficiencies of viscous dissipation, a significant number of our BH+OB star models with orbital periods 100 d can produce observable X-rays in the LMC (see also Fig. 6, for = 0.1).
@@ 4.5 Absolute number of X-ray-faint BH+OB star binaries in the LMC
Of the 126 BH+OB star binaries that are not expected to form an accretion disc around the BH, the number of X-ray-faint BH+OB star binaries depend on the efficiency of viscous heating adopted in our models. The predicted numbers range among 28…44…72 for = 0.001…0.1…0.5, respectively. Thus, a significant handful of BH+OB star binaries in the LMC are expected to produce faint yet observable X-ray emission, even for the most inefficient case of viscous heating. Moreover, if accretion disc formation is inhibited in some of the BH+OB binaries that are predicted to be X-ray-bright in this work, the high mass-accretion rates predicted in the X-ray-bright systems will make them observable as X-ray-faint sources. Hence, the number of X-ray-faint BH+OB star binaries will increase for more stringent constraints on the accretion disc formation criterion (Sect. 6.4).
@@ 5 Comparison with observations
Table 1 lists confirmed and tentative BH+OB star binaries observed in the LMC. LMC X-1 is a long-studied system with continuous X-ray emission produced from a Keplerian accretion disc (see, e.g. Pakull & Angebault 1986; Orosz et al. 2009). VFTS 243 is a recently discovered X-ray inactive BH+O star binary in the Tarantula Nebula (Shenar et al. 2022a). Shenar et al. (2022b) identified two more systems, VFTS 514 and 779, as strong candidates to host BHs that do not show strong X-ray emission. The remaining systems are SB1 binaries with unseen companions of mass above 2.35 333The most massive neutron star discovered is 2.35 (Romani et al. 2022) (Table 2 of Shenar et al. 2022b).
@@ 5.1 X-ray-bright systems
In the right panel of Fig. 3, we use the criterion for the formation of an accretion disc to show the domains where a disc is expected or not (with and ), for black hole masses corresponding to three high-mass X-ray binaries: LMC X-1 (Orosz et al. 2009), Cygnus X-1 (Orosz et al. 2011) and M33 X-7 (Orosz et al. 2007; Valsecchi et al. 2010; Ramachandran et al. 2022). The wind speed of the main sequence companion to the BH is computed at the orbital separation with a -law (Eq. 1 of Sen et al. 2021). Each coloured circle marker lies below its respective threshold for accretion disc formation. This shows that even if the OB star companion does not fill its Roche lobe, an accretion disc can form through wind capture in these three systems, from where copious amounts of X-rays can be emitted. In the grey region, the wind launching is quenched and mass accretion onto the BH can only proceed through Roche lobe overflow.
We see that the binary models cannot reproduce the position of LMC X-1 in the orbital period-wind speed plane. The mass of the BH and the O star companion is 10.911.41 and 31.793.48 respectively. However, our models at the relevant initial mass (30-50 , which can form a 10 BH), and initial orbital period (1-3 d) range merge on the main sequence after a contact phase. This may indicate the short period may detach after a brief contact phase. Investigating the physics of the contact phase is beyond the scope of this work (see, e.g. Fabry et al. 2022, 2023). Alternatively, at an orbital period of 3.9 d the black hole in LMC X-1 is deeply engulfed in the wind acceleration region of the 31.8 O star. X-rays from the BH can ionise the stellar wind decreasing its speed (Krtička & Kubát 2009). This effect is not taken into account in our simple analysis. We assume that the O star companion has a radiatively driven wind which follows the -law (as in Orosz et al. 2007).
On the other hand, our models predict a subpopulation of X-ray-bright BH+OB star binaries above 15 d with wind speeds below 800 km. The OB star companions in this population also have a low filling factor. There are no observed wind-fed HMXBs in this region. Despite the low-number statistics, this may imply two possibilities. First, binaries may not undergo stable mass transfer at low initial mass ratios and long orbital periods (see however, Schürmann & Langer 2024, where the criterion for unstable mass transfer is more relaxed). On the other hand, our criterion to determine the formation of an accretion disc may be inadequate to filter out BH+OB models where the OB star has low filling factors. We discuss this further in Sect. 6.4.

### v3

@@ 3.3 X-ray luminosity from a Keplerian accretion disk
where is the wind loss rate of the OB star, is the orbital separation, is the accretion radius of the BH, is the wind velocity of the O star, is the relative velocity of the BH with respect to the wind velocity (see equations (1)-(8) of Sen et al. 2021), and represents the self-limiting effect of the Eddington luminosity,
@@
where , and are the mass, radius and Eddington factor of the OB star companion.
3.4 X-ray luminosity without an accretion disk
When the specific angular momentum carried by the OB star wind matter is insufficient to form a Keplerian accretion disk around the BH, the wind matter can spiral into the BH following the magnetic field lines. During this in-fall, we have shown in Sect. 2.2.3 that electrons can receive a portion of the viscously dissipated energy of the in-falling matter, and emit radiation in the X-ray band. The X-ray luminosity is
where is the radiative efficiency and is the net mass accretion rate at the radius of the ISCO of the BH.
Recent simulations of hot accretion flows have revealed the presence of outflows in ADAFs, such that the mass inflow rate of the accreting gas decreases as the material approaches the event horizon of the BH (Stone et al. 1999; Yuan & Bu 2010). We assume the mass accretion rate scales with the radius from the BH by a power-law (, Stone et al. 1999; Igumenshchev et al. 2000; Stone & Pringle 2001; Yuan & Bu 2010). Hence, the relation between the net mass accretion rate at the radius of the innermost stable circular orbit and the mass accretion rate at the accretion radius is given by
@@ 3.4 X-ray luminosity without an accretion disk
Xie & Yuan (2012, Fig. 1) studied the radiative efficiency of hot accretion flows around a stellar-mass BH orbiting an OB star companion. They derive as a function of the net mass accretion rate (i.e. = (,)), for three different values of the viscous dissipation parameter (defined in Sect. 2.2.3). From their figure 1, we extract the values of the radiative efficiency as a function of the net mass accretion rate to calculate the X-ray luminosity during the BH+OB star phase in our models for viscous dissipation parameter . While we show in Sect. 2 that viscous heating of electrons can be efficient even in the case of stellar mass BHs with OB star companions, we present our results for all three values above to derive a lower limit on the number of faint X-ray sources from BH+OB star binaries for the most inefficient case of viscous heating.
@@ 4 Results
4.1 Mass accretion rate
Figure 1 shows the distribution of the mass accretion rate at the accretion radius (left panel) and the net mass accretion rate near the event horizon of the BH (right panel). Higher mass accretion rates are reached for shorter orbital period systems (Eq. (22), expressed in terms of orbital separation). The peak in the mass accretion rate distribution is and for 1.4-10 d and 10-100 d initial orbital periods respectively. The shortest-period systems can also form an accretion disk around the BH (Sen et al. 2021), denoted by the black hatching. We note however that there is a significant contribution from the 10-100 d binaries to the population of BH+OB star systems that can form an accretion disk (see Sect. 4.3).
The highest mass accretion rates () is mostly from models in the 10-100 d initial orbital period range. The shortest initial orbital period models ( d) enter a contact phase and merge on the main sequence (Menon et al. 2021). Many of the short-period binary models (2-10 d) in our grid do not survive their prior Case A and Case AB mass transfer phase (Sen et al. 2022, Fig. 1 of, Pauli et al. 2022, Appendix C of) to reach the BH+OB star phase.
The total number of short-period (1.4-10 d) binary models contributing to the distribution of mass accretion rate is much smaller than the number of models with longer initial orbital periods (10-100 d). Correspondingly, the peak in the distribution of mass accretion rate does not occur at the highest mass accretion rates but at lower mass accretion rates where more models contribute to the histogram function (Eq. (33)), despite our assumed intrinsic period distribution favouring short-period binaries (Eq. (34)).
The mass accretion rate typically decreases for models with longer orbital periods (Eq. (22)). The drop in the number of systems with mass accretion rates below is produced by our assumption that longer-period binaries are less likely to be born (Eq. (34)). Hence, the peak in the distribution of mass accretion rates arises at 10-12-10-11 yr-1, from BH+OB star binaries with intermediate orbital periods.
We assume that mass outflows reduce the mass accretion rate from the accretion radius to the BH event horizon (Eq. 32). The right panel shows that the peak in the distribution of the net mass accretion rate near the event horizon of the BH is 2 orders of magnitude lower than that at the accretion radius. This reduction in mass accretion rate results in a proportionate decrease in X-ray luminosity (Eq. 31) predicted from ADAFs around BHs in the BH+OB star binaries. We note that observations have confirmed the presence of outflows in Sgr A* (Hawley & Balbus 2002; Yuan et al. 2003; Igumenshchev et al. 2003). In case the same may not true for accretion onto stellar-mass BHs (e.g., see Fender & Gallo 2014), our predictions for the X-ray luminosity will be 2 orders of magnitude higher than presented in the next section.
@@ 4.2 X-ray luminosity
Figure 2 shows the X-ray luminosity distribution from BH+OB star binaries according to our binary evolution models due to accretion onto the BH (coloured histograms) and shocks in the wind of the OB star companion (black step histogram). In all the panels, the X-ray luminosity in the range erg s-1 originate solely from a Keplerian accretion disk (Shakura 1973) around the BHs (Eq. 25). These strong X-ray sources comprise 7.85% of the entire population of BH+OB star binaries in our grid. The orbital period distribution (see figure 6 of Langer et al. 2020) of the BH+OB star binaries peaks above 100 d, where the strong wind velocity of the OB star companion disfavours the formation of an accretion disk (Sen et al. 2021). Hence, most BH+OB star binaries in our model grid do not form an accretion disk. We call the BH+OB star binaries with X-ray luminosity above erg s-1 ‘X-ray-bright’.
We find that the contribution from binary models in the 40-90 range to the population of X-ray-bright BH+OB star binaries is negligible. We identify two reasons. Firstly, the IMF strongly disfavours the most massive binaries in the population. Secondly, the most massive models enter a contact phase up to an initial orbital period of 3 d (figures C1-C3 of Pauli et al. 2022). Hence, the shortest-period models, likely to form an accretion disk during a BH+OB phase, do not reach their BH+OB star phase.
The X-ray luminosity distribution shows a second, broader peak at luminosities in the range from erg s-1 which is associated with systems having ADAFs (Narayan & Yi 1995) around BHs that do not have Keplerian accretion disk around them (Sect. 3.4). The largest contribution comes from the binaries with initial primary masses of 20-40 . A large fraction of the primaries in the initial mass range of 10-20 have helium core mass smaller than 6.6 such that they are not expected to collapse into BHs while the contribution from the 40-90 range is suppressed by the IMF and comparatively shorter lifetimes of more massive O-star companions.
@@
The X-ray luminosity of observed Be X-ray binaries is higher than 1034 erg s-1, even for the faint persistent sources (Cheng et al. 2014). For inefficient viscous heating of electrons, the X-ray luminosity predicted from the BH+O star binaries remain below the lower limit of the X-ray luminosity from Be X-ray binaries (top and middle panels). For = 0.5 (bottom panel), only 2% of the BH+O star binaries may have X-ray luminosity between erg s-1, which is lower than the fraction of BH+O star binaries predicted to have a Keplerian accretion disk around the BH. We discuss this in more detail in Sect. 6.2.
@@ 4.3 X-ray-bright systems
The left panel of Fig. 3 shows the distribution of the Roche lobe filling factor of the OB star companion when an accretion disk can form during the BH+OB star phase of the binary models. We find a distinct bimodal distribution, partly as a natural consequence of the binary orbital periods (1.4 d-10d vs 10-100 d) that contribute to the different peaks (0.3 vs 0.9). The filling factors of models arising from 1.4-10 d models peak at 0.9. The orbital periods are short enough that the accretion disk can form only towards the end of the main sequence of the O star companions when the stellar radius increases and the wind speed decreases. On the other hand, the contribution from 10-100 d binary models comes primarily from BHs with B star companions, where the terminal velocity of the B star wind is a factor of two lower than O stars. Moreover, many binary models with small initial mass ratios merge on the main sequence at the 1.4-10 d orbital period range (e.g., figure 1 and figure F1 of Sen et al. 2022).
@@ 4.4 X-ray-faint systems
Figure 4 shows the orbital period and OB star luminosity of the BH+OB star binaries that are predicted to emit X-ray emission above 1031 erg s-1, the observational threshold of current X-ray telescopes/surveys (Crowther et al. 2022). For = 0.001, the predicted distribution of orbital periods peak near 10 d. The mass accretion rate falls inversely with the square of the orbital separation (Eq. 22), thereby making higher orbital period binaries less likely to overcome the X-ray luminosity threshold in the LMC. At the shortest orbital periods, all our binary models do not survive the mass transfer phases before the BH+OB star phase (Sect. 4.1), due to which the probability distribution decreases below 10 d.
The wind mass loss rate of the OB stars increases with luminosity (see, e.g., Kudritzki & Puls 2000; Puls et al. 2008) and mass (Vink et al. 2001a). BHs in orbit with more massive (luminous) stars at the same orbital period have a greater mass accretion rate than less massive (luminous) stars (Eq. 22), as the mass accretion rate is proportional to the wind mass-loss rate. Consequently, the luminosity of the OB star has to be higher at long orbital periods to keep the mass accretion rate onto the BH high enough to produce observable X-ray emission. As such, we see that the distribution of observable BH+OB star binaries extends towards higher luminosities at higher orbital periods.
For = 0.5 (right panel of Fig. 4), the radiative efficiency increases by 2 orders of magnitude compared to = 0.001, which leads to higher X-ray luminosity produced from the same mass accretion rate. We see a larger contribution from longer-period binaries to the distribution of BH+OB star binaries with observable X-ray emission. The increased radiative efficiency compensates for the lower mass accretion rate at longer periods. For high efficiencies of viscous dissipation, a significant number of our BH+OB star models with orbital periods 100 d can produce observable X-rays in the LMC (see also Fig. 6, for = 0.1).
@@ 4.5 Absolute number of X-ray-faint BH+OB star binaries in the LMC
Of the 126 BH+OB star binaries that are not expected to form an accretion disk around the BH, the number of X-ray-faint BH+OB star binaries depend on the efficiency of viscous heating adopted in our models. The predicted numbers range among 28…44…72 for = 0.001…0.1…0.5, respectively. Thus, a significant handful of BH+OB star binaries in the LMC are expected to produce faint yet observable X-ray emission, even for the most inefficient case of viscous heating. Moreover, if accretion disk formation is inhibited in some of the BH+OB binaries that are predicted to be X-ray-bright in this work, the high mass accretion rates predicted in the X-ray-bright systems will make them observable as X-ray-faint sources. Hence, the number of X-ray-faint BH+OB star binaries will increase for more stringent constraints on the accretion disk formation criterion (Sect. 6.4).
@@ 5 Comparison with observations
Table 1 lists confirmed and tentative BH+OB star binaries observed in the LMC. LMC X-1 is a long-studied system with continuous X-ray emission produced from a Keplerian accretion disk (see, e.g. Pakull & Angebault 1986; Orosz et al. 2009). VFTS 243 is a recently discovered X-ray inactive BH+O star binary in the Tarantula Nebula (Shenar et al. 2022a). Shenar et al. (2022b) identified two more systems, VFTS 514 and 779, as strong candidates to host BHs that do not show strong X-ray emission. The remaining systems are SB1 binaries with unseen companions of mass above 2.35 333The most massive neutron star discovered is 2.35 (Romani et al. 2022) (Table 2 of Shenar et al. 2022b).
@@ 5.1 X-ray-bright systems
In the right panel of Fig. 3, we use the criterion for the formation of an accretion disk to show the domains where a disk is expected or not (with and ), for black hole masses corresponding to three high-mass X-ray binaries: LMC X-1 (Orosz et al. 2009), Cygnus X-1 (Orosz et al. 2011) and M33 X-7 (Orosz et al. 2007; Valsecchi et al. 2010; Ramachandran et al. 2022). The wind speed of the main sequence companion to the BH is computed at the orbital separation with a -law (equation 1 of Sen et al. 2021). Each coloured circle marker lies below its respective threshold for accretion disk formation. This shows that even if the OB star companion does not fill its Roche lobe, an accretion disk can form through wind capture in these three systems, from where copious amounts of X-rays can be emitted. In the grey region, the wind launching is quenched and mass accretion onto the BH can only proceed through Roche lobe overflow.
We see that the binary models cannot reproduce the position of LMC X-1 in the orbital period-wind speed plane. The mass of the BH and the O star companion is 10.911.41 and 31.793.48 respectively. However, our models at the relevant initial mass (30-50 , that can form a 10 BH), and initial orbital period (1-3 d) range merge on the main sequence after a contact phase. This may indicate the short period may detach after a brief contact phase. Investigating the physics of the contact phase is beyond the scope of this work (see, e.g. Fabry et al. 2022, 2023). Alternatively, at an orbital period of 3.9 d the black hole in LMC X-1 is deeply engulfed in the wind acceleration region of the 31.8 O star. X-rays from the BH can ionize the stellar wind decreasing its speed (Krtička & Kubát 2009). This effect is not taken into account in our simple analysis. We assume that the O star companion has a radiatively driven wind which follows the -law (as in Orosz et al. 2007).
On the other hand, our models predict a sub-population of X-ray-bright BH+OB star binaries above 15 d with wind speeds below 800 km. The OB star companions in this population also have a low filling factor. There are no observed wind-fed HMXBs in this region. Despite the low-number statistics, this may imply two possibilities. First, binaries may not undergo stable mass transfer at low initial mass ratios and long orbital periods (see however, Schürmann & Langer 2024, where the criterion for unstable mass transfer is more relaxed). On the other hand, our criterion to determine the formation of an accretion disk may be inadequate to filter out BH+OB models where the OB star has low filling factors. We discuss this further in Sect. 6.4.